Shift report 20000701 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sat Jul 01 2000 - 23:58:29 EDT

  • Next message: http account: "Shift report 20000702 00:00-08:00"

    ------------------------------------------------------------
    Date: Sat Jul  1 23:58:29 2000, Shift: 16:00-24:00
    Supervisor: R. Debbe, with: B. Samset +
    Log book#2 pages 106 - 110
    
    Shift Summary:
    ------------------------------------------------------------
    Two stores during this shift: Took data from 18:00 to 20:00
    A quench link event killed the first store.
    Second store started at 21:30 and is still going around the pipes.
    We took data with trigger 6 to check timing of C1 and tiles.
    
    D1 and D2 magnet have been turned on at the begining of the shift, they have 1/4 of max field.
    Bjoern exercised the vertex finder on data taken during this shift.
    He reconstructed tracks in MTPC1.
    
    T2 is very quiet. C1 is detecting light well timed with the 
    trigger, T2 should have tracks going thru it.
    
    We are now taking data with triggers 4,6,8
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sun Jul 02 2000 - 00:03:03 EDT